Measuring the Cost of Living (Mobile-Enabled) Graded Assignment | Read Chapter 11| B ack to Due Sunday 02.25.18 at 11:45 F Keep the Highest: /3 Attempts 5. Interest, inflation, and purchasing power Suppose Luca is a cinephile and buys only movie tickets. Lucia deposits $1,000 in a bank account that pays an annual nominal interest rate of 5% Assume this interest rate is fixed-that is, it wont change over time. At the time of her deposit, a movie ticket is priced at $10.00. Initially, the purchasing power of Lucias $1,000 deposit is movie tickets. For each of the annual inflation rates given in the following table, first determine the new price of a movie ticket, assuming it rises at the rate of inflation. Then enter the corresponding purchasing power of Lucias deposit after one year in the first row of the table for each inflation rate. Finally, enter the value for the real interest rate at each of the given inflation rates. Hint: Round your answers in the first row down to the nearest movie ticket. For example, if you find that the deposit will cover 20.7 movie tickets, you would round the purchasing power down to 20 movie tickets under the assumption that Lucia will not buy seven-tenths of a movie ticket. Annual Inflation Rate 596 0% 10% Number of Tickets Lucia Can Purchase after One Year Real Interest Rate When the rate of inflation is equal to the interest rate on Lucias deposit, the purchasing power of her deposit of the year. over the course Grade It Now Save & Continue
Answer #1

Answer :As initial deposit with Lucia = $1000

Price of the movie ticket = $10

Nominal interest rate = 5%

Number of movie ticket = Deposits/ Price of the ticket

Number of movie ticket = 1000/10

Number of movie ticket =100

Initially the purchasing power of Lucia $1000 deposit is 100 tickets.

Lucia amount hold after one year deposit is $1050

S.I = Principal*Rate of interest *Time

S.I = $1000*5/100*1

S.I = $50

Amount = P+I = $1000+$50 = $1050

Annual inflation rate 0% 5% 10%
Number of ticket purchased after one year 1050/10 = 105 1050/10.5=100 1050/11=95
Amount after one year $1050 $1050 $1050

Real interest rate

r=nominal -inflation

r=5%-0% = 5% r=5%-5% = 0

r=5%-10%

=-5%

When rate of inflation is equal to the interest rate of lucia deposit the purchasing power of her deposit is equal to its purchasing power over the course of the year. It means that before one year Lucia should purchase 100 tickets and after one year with same inflation rate as interest rate 100 tickets when inflation rate is 5%.